# Object-Centric Latent Action Learning

### v2

@@ Abstract
Leveraging vast amounts of unlabeled internet video data for embodied AI is currently bottlenecked by the lack of action labels and the presence of action-correlated visual distractors. Although recent latent action policy optimization (LAPO) has shown promise in inferring proxy-action labels from visual observations, its performance degrades significantly when distractors are present. To address this limitation, we propose a novel object-centric latent action learning framework that centers on objects rather than pixels. We leverage self-supervised object-centric pretraining to disentangle action-related and distracting dynamics. This allows LAPO to focus on task-relevant interactions, resulting in more robust proxy-action labels, enabling better imitation learning and efficient adaptation of the agent with just a few action-labeled trajectories. We evaluated our method in eight visually complex tasks across the Distracting Control Suite (DCS) and Distracting MetaWorld (DMW). Our results show that object-centric pretraining mitigates the negative effects of distractors by 50%, as measured by downstream task performance: average return (DCS) and success rate (DMW).
@@ 1 Introduction
Figure 1: Main Results. We achieve significant downstream performance improvements across all environments and budgets of action-labeled trajectories considered. We normalize returns and success rates by the performance of the BC agent trained on the all ground-truth action labels. LAPO-vanilla uses data without distractors, representing the upper bound performance. Results averaged over three random seeds. See Section 4 for additional details.
@@
In recent years, the scaling of model and data sizes has led to the creation of powerful and general foundation models (Bommasani et al., 2021) that have enabled many breakthroughs in understanding and generation of natural language (Achiam et al., 2023; Brown et al., 2020) and images (Dehghani et al., 2023; Radford et al., 2021). In contrast, the field of embodied AI has generally remained behind in terms of generalization and emergent abilities, being mostly limited by the lack of diverse high-quality data for pre-training (Guruprasad et al., 2024; Lin et al., 2024). The vast amount of video data on the Internet, covering a wide variety of human-related activities, can potentially fulfill the current data needs for training generalist agents (McCarthy et al., 2024). Unfortunately, such videos cannot be used directly due to lack of explicit action labels, which are essential for imitation learning and reinforcement learning algorithms.
@@
In order to compensate for the lack of action labels, approaches based on Latent Action Models (LAMs) (Schmidt & Jiang, 2024; Ye et al., 2024; Cui et al., 2024; Bruce et al., 2024), aim to infer latent actions from the sequence of visual observations. Such latent actions can later be used for behavioral cloning from large unlabeled datasets. However, a major limitation of current LAM approaches, is their susceptibility to action-correlated distractors, such as dynamic backgrounds, that falsely correlate with agent actions, and may lead models to overfit to non-causal patterns (Wang et al., 2024; Misra et al., 2024b; McCarthy et al., 2024; Nikulin et al., 2025). Existing methods, such as Latent Action Pretraining (LAPA) (Ye et al., 2024; Chen et al., 2024), often assume curated, distractor-free datasets or rely on costly annotations, severely limiting their scalability and applicability in realistic settings.
@@
To address this bottleneck, we propose object-centric latent action learning, a framework that leverages object-centric representations to isolate task-relevant entities from distracting visual noise. Object-centric models decompose scenes into discrete, interpretable object slots (Locatello et al., 2020; Dittadi et al., 2022) using self-supervised mechanisms that group pixels into coherent entities based on feature similarity. These slots encode spatio-temporal properties (Zadaianchuk et al., 2023) effectively disentangling complex visual inputs into semantically meaningful units.
@@
We evaluated our approach in continuous control environments with dynamic distractors, such as Distracting Control Suite (DCS) (Stone et al., 2021) and previously unexplored Distracting Metaworld (DMW) (Yu et al., 2019), showing that self-supervised object-centric pretraining, built on VideoSAUR (Zadaianchuk et al., 2023), significantly improves robustness compared to standard LAM baselines. Our results suggest that object decomposition serves as an effective inductive bias for latent action learning, especially under realistic, noisy conditions. In addition, we propose a simple way to automatically choose slots relevant to control and analyze its effectiveness. We summarize our main results on Figure 1.
Figure 2: General overview of the our method pre-training pipeline. Object-Centric Pretraining: We decompose video sequences into interpretable object slots. A linear probe trained on slot representations automatically selects task-relevant slots by identifying those most predictive of actions. Latent Action Learning: We train a latent action model based on LAPO, learning inverse and forward dynamics in the slot space. Behavior Cloning and Fine-tuning: A behavior cloning agent is trained on the inferred latent actions. The resulting policy is then fine-tuned using a small number of trajectories with ground-truth actions ($\leq 2.5$% of total data), enabling strong downstream performance with minimal supervision.
@@ 2 Preliminaries
In our experiments, we use environments with partial observability (POMDP) (Åström, 1965; Kaelbling et al., 1998). A POMDP is defined by the tuple $(\mathcal{S},\mathcal{A},T,R,\Omega,\gamma)$, where: $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $T(s_{t+1}\mid s_{t},a_{t})$ defines the transition dynamics, $R(s_{t},a_{t},s_{t+1})$ is the reward function, $\Omega(o_{t}\mid s_{t})$ is the observation model, $\gamma\in[0,1)$ is the discount factor. At each time step $t$, the agent receives an observation, $o_{t}\sim\Omega(\cdot\mid s_{t})$, and selects an action $a_{t}\sim\pi(\cdot\mid o_{t})$ to maximize its expected discounted return: $\mathbb{E}\left[\sum_{k=t}^{\infty}\gamma^{k-t}r_{k}\right]$ The agent transitions to the next state $s_{t+1}\sim T(\cdot\mid s_{t},a_{t})$ and receives a reward $r_{t}=R(s_{t},a_{t},s_{t+1})$.
We consider the setting in which we first learn a policy $\pi$ from offline demonstration data, then fine-tune it through online interaction introduced by Schmidt & Jiang (2024). Suppose we are given a dataset $\mathcal{D}$ of expert trajectories in the form of sequences of observations and actions:
@@
When such labeled data is available, Behavior Cloning (BC) (Pomerleau, 1988) can be used to train a policy $\pi_{\text{BC}}(a_{t}\mid o_{t})$ through supervised learning. However, when using real-world video data the action labels are missing, and we only observe sequences of the form:
@@
Edwards et al. (2019); Schmidt & Jiang (2024) proposed to convert classic BC into a pseudo-BC task by inferring latent actions $\mathbf{z}_{t}$ that explain each observed transition $(o_{t},o_{t+1})$. These inferred actions are then used to train a latent action policy $\tilde{\pi}(z_{t}\mid o_{t})$, bypassing the need for ground-truth action labels. To obtain latent actions, LAM trains an inverse-dynamics model (IDM) $f_{\mathrm{IDM}}(\boldsymbol{z}_{t}|\boldsymbol{o}_{t},\boldsymbol{o}_{t+1})$, which infers latent actions from state transitions. To ensure that the IDM learns meaningful representations, it is trained jointly with a forward-dynamics model (FDM): $f_{\mathrm{FDM}}(\boldsymbol{o}_{t+1}|\boldsymbol{o}_{t},\boldsymbol{z}_{t})$, which predicts the next observation from the current observation and the inferred action. The models are optimized to perform next observation prediction:
@@
Without distractors LAM is able to recover ground-truth actions (Schmidt & Jiang, 2024; Bruce et al., 2024), however, with distractions it does not work (Nikulin et al., 2025).
@@
Real-world videos inherently contain action-correlated distractors: environmental dynamics (e.g. moving backgrounds, camera jitter) that spuriously correlate with agent actions. However, existing learning from observation (LfO) methods lack mechanisms to disentangle distractors (Efroni et al., 2022; Misra et al., 2024a), leading to overfitting in noisy settings. Recently, Nikulin et al. (2025) empirically demonstrated that LAPO-based methods suffer significant performance degradation when trained on data that contain distractors and suggested reusing available action labels to provide supervision during LAM pre-training. However, it is not always possible to provide such supervision, since action labels do not exist in principle in some domains (e.g. YouTube videos). On the other hand, object-centric decomposition can be applied in a meaningful way to any real-world video data, greatly expanding the applicability of latent action learning.
@@
Object-centric learning (OCL) aims to decompose complex visual scenes into structured, interpretable representations by isolating individual entities over time. These entities are typically encoded as a set of $K$ slot vectors $\mathcal{S}_{t}=\{\boldsymbol{s}_{t}^{(1)},\boldsymbol{s}_{t}^{(2)},...,\boldsymbol{s}_{t}^{(K)}\},$ representing a distinct entity or region within the scene and capture object-specific properties such as position, shape, color, and motion. A key strength of OCL models lies in their ability to encode the causal structure of the environment. By isolating coherent, temporally consistent entities, they provide a natural defense against distractors in real-world video data. For example, VideoSAUR (Zadaianchuk et al., 2023), a self-supervised video decomposition model, has been shown to effectively ignore spurious correlations introduced by moving backgrounds or viewpoint changes, focusing instead on task-relevant objects.
Although object slots provide disentangled and semantically meaningful representations, identifying which slots correspond to task-relevant entities remains challenging and necessitates additional mechanisms to identify and track relevant entities without manual supervision. Also, most OCL frameworks, including VideoSAUR, require the number of slots $K$ to be specified a priori, which limits flexibility in modeling scenes with varying numbers of objects and can lead to either under-representation or fragmentation when $K$ is mismatched with scene complexity.
@@ 3 Method
We employ the VideoSAUR model (Zadaianchuk et al., 2023) to decompose input video frames into spatio-temporal object slots. We initially experimented with the STEVE model (Singh et al., 2022), a widely used and promising approach, but found that it failed to consistently isolate entities such as the hopper in our tasks (illustrated compared to VideoSAUR in Figure 25).
VideoSaur’s self-supervised architecture isolates individual entities within a scene, providing structured representations that are less susceptible to background noise and incidental motion. At the end of this step, we obtain an encoder, that directly maps a trajectory of observations $\tau_{o}=({o}_{1},...,{o}_{T})$ to a corresponding trajectory of slots $\tau_{s}=({s}_{1},...,{s}_{T})$. VideoSAUR encodes each observation ${o}_{t}$ in a fixed number $K$ of slot vectors ${s}_{t}^{(k)}\in\mathbb{R}^{d}$, where the parameter $K$ is set prior to training.
Due to its transformer-based decoder, VideoSAUR allows each slot ${s}_{t}^{(k)}$ to be projected back into the original observation space using attention maps as alpha masks. This enables us to visualize and interpret the spatial support of each slot in the image. We denote these by ${m}_{t}^{(k)}$ and refer to them as object masks throughout the paper (see Figures 3, 25 and 26 for visualizations).
To improve consistency between training runs and reduce slot permutation issues, we employ fixed initialization of slots (see details in Appendix C), which ensures that the same slot index corresponds to a similar semantic object across different episodes.
Slot Selection.
In our setting, control-relevant objects typically include the main agent (e.g. cheetah in "cheetah-run" from DCS, robotic arm and gripper from DWM) and some objects (hammer, cubes, etc. in DMW), depending on the environment and the number of slots $K$, these entities can occupy one or more slots. To identify which slots correspond to the agent’s interactions, we propose an automated method to select relevant slots based on their relationship to inferred actions, inspired by Alain & Bengio (2016). A linear regressor is trained to predict ground-truth actions from fixed slot encodings on a small number of labeled trajectories. We apply Principal Component Analysis (PCA) to slot encodings beforehand to reduce dimensionality and noise.
To ensure robustness and reduce variance in performance estimates, we use 5-fold cross-validation and report the average mean squared error (MSE) on the test data, as the final evaluation metric. Slots with lower MSE are considered to be more predictive of the true action and thus more likely to represent control-relevant entities. We select a subset of top-$N$ slots with the lowest prediction error as our final set of relevant slots $\mathcal{S}_{t}^{\star}=\{\boldsymbol{s}_{t}^{(k)}\mid k\in\mathcal{K}^{\star}\}$, where $\mathcal{K}^{\star}$ denotes the indices of selected slots. This selection process is performed once per dataset due to the fixed initialization of slots, ensuring consistent interpretation across episodes.
@@
Utilizing the selected object-centric representations, we train a latent action model inspired by LAPO (Schmidt & Jiang, 2024) under the space of object-centric representations. The inverse-dynamics model $\boldsymbol{z}_{t}\sim f^{s}_{\mathrm{IDM}}(\cdot|\boldsymbol{s}_{t},\boldsymbol{a}_{t+1})$ and the forward-dynamics model $\boldsymbol{\hat{s}}_{t+q}\sim f^{s}_{\mathrm{FDM}}(\cdot|\boldsymbol{s}_{t},\boldsymbol{z}_{t})$ are trained to reconstruct the trajectory of slots ${||\boldsymbol{\hat{s}}_{t+1}-\boldsymbol{s}_{t+1}||}^{2}$ (or masks). We denote this as LAPO-slots and, accordingly, LAPO-masks for masks representations. Thus, LAPO-slots reconstruct next observations in latent space, while LAPO-masks reconstruct images, which were filtered based on selected slots (see Figures 3 and 26).
@@
The inferred latent actions serve as proxies for actual action labels. We train a behavior cloning (BC) agent to predict these latent actions, using the same dataset as for latent action learning. To evaluate the pre-training effectiveness, as a final stage, we fine-tune the resulting agent on a limited set of trajectories with ground-truth action labels (no more than 2.5% of total data), in line with (Schmidt & Jiang, 2024; Ye et al., 2024; Nikulin et al., 2025). The BC agent scores for different numbers of labeled trajectories are presented on Figure 4.

### v3

@@ Abstract
Leveraging vast amounts of unlabeled internet video data for embodied AI is currently bottlenecked by the lack of action labels and the presence of action-correlated visual distractors. Although recent latent action policy optimization (LAPO) has shown promise in inferring proxy action labels from visual observations, its performance degrades significantly when distractors are present. To address this limitation, we propose a novel object-centric latent action learning framework that centers on objects rather than pixels. We leverage self-supervised object-centric pretraining to disentangle the movement of the agent and distracting background dynamics. This allows LAPO to focus on task-relevant interactions, resulting in more robust proxy-action labels, enabling better imitation learning and efficient adaptation of the agent with just a few action-labeled trajectories. We evaluated our method in eight visually complex tasks across the Distracting Control Suite (DCS) and Distracting MetaWorld (DMW). Our results show that object-centric pretraining mitigates the negative effects of distractors by 50%, as measured by downstream task performance: average return (DCS) and success rate (DMW).
Codehttps://github.com/dunnolab/object-centric-lapo
@@ 1 Introduction
In recent years, the scaling of model and data sizes has led to the creation of powerful and general foundation models (Bommasani et al. 2021) that have enabled many breakthroughs in understanding and generation of natural language (Achiam et al. 2023; Brown et al. 2020) and images (Dehghani et al. 2023; Radford et al. 2021). In contrast, the field of embodied AI has generally remained behind in terms of generalization and emergent abilities, being mostly limited by the lack of diverse high-quality data for pre-training (Guruprasad et al. 2024; Lin et al. 2024). The vast amount of video data on the Internet, covering a wide variety of human-related activities, can potentially fulfill the current data needs for training generalist agents (McCarthy et al. 2024). Unfortunately, such videos cannot be used directly due to lack of explicit action labels, which are essential for imitation learning and reinforcement learning algorithms.
Figure 1: Main Results. Our methods, LAPO-slots and LAPO-masks, leverage object-centric representations to significantly improve latent action learning under visual distractions. Compared to baseline (LAPO (Schmidt and Jiang 2024), trained on raw videos with distraction), our approaches reduce the performance gap toward the clean-data upper bound (LAPO-clean, trained on clean videos without distractions) by 50%. Distractions include camera movements, color variations, and dynamic background videos. Downstream performance is normalized to a behavior cloning agent trained on all ground-truth action labels. Results averaged over three random seeds. See Section˜4 for details.
@@
In order to compensate for the lack of action labels, approaches based on Latent Action Models (LAMs) (Schmidt and Jiang 2024; Ye et al. 2024; Cui et al. 2024; Bruce et al. 2024), aim to infer latent actions from the sequence of visual observations. Such latent actions can later be used for behavioral cloning from large unlabeled datasets. However, a major limitation of current LAM approaches, is their susceptibility to action-correlated distractors, such as dynamic backgrounds, that falsely correlate with agent actions, and may lead models to overfit to non-causal patterns (Wang et al. 2024; Misra et al. 2024b; McCarthy et al. 2024; Nikulin et al. 2025). Existing methods, such as Latent Action Pretraining (LAPA) (Ye et al. 2024; Chen et al. 2024), often assume curated, distractor-free datasets or rely on costly annotations, severely limiting their scalability and applicability in realistic settings.
Figure 2: General overview of our method pre-training pipeline. Object-Centric Pretraining: We decompose video sequences into interpretable object slots. A linear probe trained on slot representations automatically selects task-relevant slots by identifying those most predictive of actions. Latent Action Learning: We train a latent action model based on LAPO, learning inverse and forward dynamics in the slot space. Behavior Cloning and Fine-tuning: A behavior cloning agent is trained on the inferred latent actions. The resulting policy is then fine-tuned using a small number of trajectories with ground-truth actions ($\leq 2.5$% of total data), enabling strong downstream performance with minimal supervision.
@@
To address this bottleneck, we propose object-centric latent action learning, a framework that leverages object-centric representations to isolate task-relevant entities from distracting visual noise. Object-centric models decompose scenes into discrete, interpretable object slots (Locatello et al. 2020; Dittadi et al. 2022) using self-supervised mechanisms that group pixels into coherent entities based on feature similarity. These slots encode spatio-temporal properties (Zadaianchuk et al. 2023) effectively disentangling complex visual inputs into semantically meaningful units.
@@
We evaluated our approach in continuous control environments with dynamic distractors, such as Distracting Control Suite (DCS) (Stone et al. 2021) and previously unexplored Distracting Metaworld (DMW) (Yu et al. 2019), showing that self-supervised object-centric pretraining, built on VideoSAUR (Zadaianchuk et al. 2023), significantly improves robustness compared to standard LAM baselines. Our results suggest that object decomposition serves as an effective inductive bias for latent action learning, especially under realistic, noisy conditions. In addition, we propose a simple way to automatically choose slots relevant to control and analyze its effectiveness. We summarize our main results on Figure˜1.
@@ 2 Preliminaries
In our experiments, we use environments with partial observability (POMDP) (Åström 1965; Kaelbling et al. 1998). A POMDP is defined by the tuple $(\mathcal{S},\mathcal{A},T,R,\Omega,\gamma)$, where: $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $T(s_{t+1}\mid s_{t},a_{t})$ defines the transition dynamics, $R(s_{t},a_{t},s_{t+1})$ is the reward function, $\Omega(o_{t}\mid s_{t})$ is the observation model, $\gamma\in[0,1)$ is the discount factor. At each time step $t$, the agent receives an observation, $o_{t}\sim\Omega(\cdot\mid s_{t})$, and selects an action $a_{t}\sim\pi(\cdot\mid o_{t})$ to maximize its expected discounted return: $\mathbb{E}\left[\sum_{k=t}^{\infty}\gamma^{k-t}r_{k}\right]$ The agent transitions to the next state $s_{t+1}\sim T(\cdot\mid s_{t},a_{t})$ and receives a reward $r_{t}=R(s_{t},a_{t},s_{t+1})$.
We consider the setting in which we first learn a policy $\pi$ from offline actionless demonstration data, then fine-tune it on real actions, introduced by (Schmidt and Jiang 2024). Suppose we are given a dataset $\mathcal{D}$ of expert trajectories in the form of sequences of observations and actions:
@@
When such labeled data is available, Behavior Cloning (BC) (Pomerleau 1988) can be used to train a policy $\pi_{\text{BC}}(a_{t}\mid o_{t})$ through supervised learning. However, when using real-world video data the action labels are missing, and we only observe sequences of the form:
@@
Edwards et al. (2019); Schmidt and Jiang (2024) proposed to convert classic BC into a pseudo-BC task by inferring latent actions $\mathbf{z}_{t}$ that explain each observed transition $(o_{t},o_{t+1})$. These inferred actions are then used to train a latent action policy $\tilde{\pi}(z_{t}\mid o_{t})$, bypassing the need for ground-truth action labels. To obtain latent actions, LAM trains an inverse-dynamics model (IDM) $f_{\mathrm{IDM}}(\boldsymbol{z}_{t}|\boldsymbol{o}_{t},\boldsymbol{o}_{t+1})$, which infers latent actions from state transitions. To ensure that the IDM learns meaningful representations, it is trained jointly with a forward-dynamics model (FDM): $f_{\mathrm{FDM}}(\boldsymbol{o}_{t+1}|\boldsymbol{o}_{t},\boldsymbol{z}_{t})$, which predicts the next observation from the current observation and the inferred action. The models are optimized to perform next observation prediction:
@@
Without distractors LAM is able to recover ground-truth actions (Schmidt and Jiang 2024; Bruce et al. 2024), however, with distractions it does not work (Nikulin et al. 2025).
@@
Real-world videos inherently contain action-correlated distractors: environmental dynamics (e.g. moving backgrounds, camera jitter) that spuriously correlate with agent actions. However, existing learning from observation methods lack mechanisms to disentangle distractors (Efroni et al. 2022; Misra et al. 2024a), leading to overfitting in noisy settings. Recently, Nikulin et al. (2025) empirically demonstrated that LAPO-based methods suffer significant performance degradation when trained on data that contain distractors and suggested reusing available action labels to provide supervision during LAM pre-training. However, it is not always possible to provide such supervision, since action labels do not exist in principle in some domains (e.g. YouTube videos). On the other hand, object-centric decomposition can be applied in a meaningful way to any real-world video data, greatly expanding the applicability of latent action learning.
@@
Object-centric learning (OCL) aims to decompose complex visual scenes into structured, interpretable representations by isolating individual entities over time. These entities are typically encoded as a set of $K$ slot vectors $\mathcal{S}_{t}=\{\boldsymbol{s}_{t}^{(1)},\boldsymbol{s}_{t}^{(2)},...,\boldsymbol{s}_{t}^{(K)}\},$ representing a distinct entity or region within the scene and capture object-specific properties such as position, shape, color, and motion. A key strength of OCL models lies in their ability to encode the causal structure of the environment. By isolating coherent, temporally consistent entities, they provide a natural defense against distractors in real-world video data. For example, VideoSAUR (Zadaianchuk et al. 2023), a self-supervised video decomposition model, has been shown to effectively ignore spurious correlations introduced by moving backgrounds or viewpoint changes, focusing instead on task-relevant objects.
Although object slots provide disentangled and semantically meaningful representations, identifying which slots correspond to task-relevant entities remains challenging and necessitates additional mechanisms to identify and track relevant entities without manual supervision. Also, most OCL frameworks, require the number of slots $K$ to be specified a priori, which limits flexibility in modeling scenes with varying numbers of objects and can lead to either under-representation or fragmentation when $K$ is mismatched with scene complexity.
@@ 3 Method
We employ the VideoSAUR model (Zadaianchuk et al. 2023) to decompose input video frames into spatio-temporal object slots. We initially experimented with the STEVE model (Singh et al. 2022), a widely used and promising approach, but found that it failed to consistently isolate entities such as the hopper in our tasks (see Supplementary Material). VideoSaur’s self-supervised architecture isolates individual entities within a scene, providing structured representations that are less susceptible to background noise and incidental motion. At the end of this step, we obtain an encoder, that directly maps a trajectory of observations $\tau_{o}=({o}_{1},...,{o}_{T})$ to a corresponding trajectory of slots $\tau_{s}=({s}_{1},...,{s}_{T})$. VideoSAUR encodes each observation ${o}_{t}$ in a fixed number $K$ of slot vectors ${s}_{t}^{(k)}\in\mathbb{R}^{d}$, where the parameter $K$ is set prior to training. Due to its transformer-based decoder, VideoSAUR allows each slot ${s}_{t}^{(k)}$ to be projected back into the original observation space using attention maps as alpha masks. This enables us to visualize and interpret the spatial support of each slot in the image. We denote these by ${m}_{t}^{(k)}$ and refer to them as object masks throughout the paper (see Figure˜3 and Supplementary Material for visualizations). To improve consistency between training runs and reduce slot permutation issues, we employ fixed initialization of slots (see details in Supplementary Material), which ensures that the same slot index corresponds to a similar semantic object across different episodes.
Environment & Task & LAPO LAPO-clean LAPO-masks LAPO-slots
Metric mean over tasks baseline upper-bound ours +gain% ours +gain%
DCS-Hard (Normalized Return) cheetah-run 0.24 ± 0.02 0.76 ± 0.04 0.41 ± 0.03 +32% 0.55 ± 0.04 +58%
hopper-hop 0.03 ± 0.01 0.27 ± 0.03 0.08 ± 0.01 +20% 0.15 ± 0.02 +50%
walker-run 0.04 ± 0.01 0.32 ± 0.07 0.06 ± 0.01 +6% 0.12 ± 0.01 +27%
humanoid-walk 0.02 ± 0.01 0.06 ± 0.01 0.04 ± 0.02 +47% 0.06 ± 0.01 +105%
mean 0.08 ± 0.01 0.35 ± 0.04 0.15 ± 0.02 +26% 0.22 ± 0.02 +52%
DCS (Normalized Return) cheetah-run 0.34 ± 0.06 0.76 ± 0.04 0.54 ± 0.09 +48% 0.50 ± 0.07 +38%
hopper-hop 0.05 ± 0.01 0.27 ± 0.03 0.15 ± 0.01 +46% 0.19 ± 0.02 +64%
walker-run 0.10 ± 0.01 0.32 ± 0.07 0.21 ± 0.05 +52% 0.18 ± 0.03 +36%
humanoid-walk 0.02 ± 0.01 0.06 ± 0.01 0.05 ± 0.02 +67% 0.09 ± 0.02 +174%
mean 0.13 ± 0.02 0.35 ± 0.04 0.24 ± 0.04 +50% 0.24 ± 0.03 +50%
DMW (Normalized Success Rate) hammer 0.75 ± 0.07 0.98 ± 0.01 0.96 ± 0.01 +91% 0.99 ± 0.02 +102%
bin-picking 0.18 ± 0.08 0.74 ± 0.10 0.49 ± 0.10 +56% 0.33 ± 0.08 +27%
basketball 0.17 ± 0.03 0.51 ± 0.07 0.34 ± 0.09 +50% 0.37 ± 0.09 +58%
soccer 0.14 ± 0.06 0.36 ± 0.08 0.21 ± 0.08 +34% 0.23 ± 0.06 +41%
mean 0.31 ± 0.06 0.65 ± 0.06 0.50 ± 0.07 +55% 0.48 ± 0.06 +50%
Table 1: Object-centric pretraining substantially improves robustness to visual distractors, where standard imitation learning from observation methods like LAPO struggle. In environments with visual distractors: dynamic background videos (DCS and DMW) and additional camera motion and color variation (DCS-Hard), LAPO’s performance is significantly degraded compared to the clean-data upper bound. We report normalized evaluation returns (DCS) and success rates (DMW) for a behavior cloning (BC) agent trained on latent actions, where a score of 1.0 corresponds to a BC agent trained on the full set of ground-truth actions. Each value averages performance across varying amounts of fine-tuning data and three random seeds. The relative performance gain is defined as: $\frac{\text{LAPO-slots/masks}\,\,-\,\,\text{LAPO}}{\text{LAPO-clean}\,\,-\,\,\text{LAPO}}\times 100\%$. This measures the fraction of the recoverable performance gap between the distracted baseline (LAPO) and the clean-data performance (LAPO-clean) that our method recovers. LAPO-slots and LAPO-masks achieve large improvements over LAPO, closing up to $\sim 50\%$ of the gap, demonstrating that object-centric representations effectively mitigate the negative effects of visual distractions.
Slot Selection via Linear Action Probe.
In our setting, control-relevant entities typically include the main agent (e.g., the cheetah in DCS tasks or the robotic arm and gripper in DMW) and task-critical objects (e.g., hammer, ball, or cubes). Depending on the environment and the number of slots $K$, these entities may be encoded across one or more slots. To automatically identify which slots capture action-relevant information, we adopt a probing approach inspired by Alain and Bengio (2016), measuring how well individual slot representations predict ground-truth actions.
We first apply Principal Component Analysis (PCA) to slot encodings computed from a small set of labeled trajectories, reducing dimensionality and noise. Then, a linear regressor is trained on resulting representations to predict the true action. To ensure robustness, we evaluate performance using 5-fold cross-validation and report the average test Mean Squared Error (MSE), which we refer to as the Linear Action Probe score. Lower MSE indicates stronger action predictivity and thus higher relevance to control.
We select the most predictive slots as the final set of relevant slots $\mathcal{S}^{\star}=\{\boldsymbol{s}^{(k)}\mid k\in\mathcal{K}^{\star}\}$, where $\mathcal{K}^{\star}$ denotes the selected slot indices. This selection is performed once per dataset after OC pretraining, leveraging the fixed slot initialization, ensuring stable and consistent slot interpretation across episodes.
@@
Utilizing the selected object-centric representations, we train a latent action model inspired by LAPO (Schmidt and Jiang 2024).
[…]
LAPO-masks operates in pixel space. It first creates a filtered image by applying the object masks from the selected slots to the input frame. The dynamics models are then trained to reconstruct this filtered image at the next timestep. Visualizations of the filtered images can be found in Figures˜3 and 5(b) and the Supplementary Material.
@@
The inferred latent actions serve as proxies for actual action labels. We train a behavior cloning (BC) agent to predict these latent actions, using the same dataset as for latent action learning. To evaluate the pre-training effectiveness, as a final stage, we fine-tune the resulting agent on a limited set of trajectories with ground-truth action labels (no more than 2.5% of total data), in line with (Schmidt and Jiang 2024; Ye et al. 2024; Nikulin et al. 2025). The finetuned BC agent scores for different numbers of labeled trajectories are presented on Figure˜4.